OpenStudy (anonymous):

For the following number, give the opposite, the reciprocal, and the absolute value. x

OpenStudy (compassionate):

\[-x, and \frac{ x }{ 1 }, and \left| x \right|\] Negative version, every x is considered to have one already over it, so switch, and the absolute value is just switch the signs.

OpenStudy (compassionate):

Actually, it's \[\frac{ 1 }{ x }\]

OpenStudy (anonymous):

The opposite is the sign of the number changed, if it is 3 then -3 is the opposite. The reciprocal is 1 divided by the number, for 3 it is 1/3. Absolute value is represented by |x|, it means that we write the positive alphanumeric value of the given number, for example |3| = 3, and |-3| = -(-3) = 3. The absolute value is an operation on the function/number.